# Personalized Lip Reading: Adapting to Your Unique Lip Movements with Vision and Language

### v1

@@ Methods
Preprocessing. By following the preprocessing (Ma et al. 2023), we crop the mouth region from talking face videos into a size of 96 pixels in both width and height. For this purpose, we employ RetinaFace (Deng et al. 2020) to detect facial landmarks and subsequently crop the mouth region based on these landmarks. Following the data augmentations described in prior research (Shi et al. 2022), we further process these cropped images by horizontally flipping them and randomly cropping them to 88 pixels in both width and height to enhance model robustness and prevent overfitting.
Architecture The design of our visual encoder incorporates a conformer architecture, configured with an embedding dimension of 768, 12 attention heads, a feed-forward dimension of 3072, and 12 blocks. The configuration of the transformer decoder, which is used for pre-training the visual encoder, includes an embedding dimension of 768, 8 attention heads, a feed-forward dimension of 3072, and 9 blocks. We employ the LLaMA3-8B (Touvron et al. 2023) model as an LLM for our proposed method.
Baseline Lip Reading Model We initialize the visual encoder with weights from a pre-trained lip reading model trained on the LRS3 dataset (Ma et al. 2023). The visual encoder is fine-tuned alongside a transformer decoder from scratch on the Baseline part of the VoxLRS-SA dataset. During the first stage of training, we employ a tri-stage learning rate with a peak learning rate of 1e-3, warmup steps of 10K, decay steps of 20K, and utilize 8 NVIDIA RTX 3090 GPUs with a maximum frames count of 1800 and gradient accumulation set to 4. In a further step of our experiment, we replace the transformer decoder with the frozen LLaMA3-8B model, continuing the training on the same datasets. In this second stage of training, we adopt a cosine learning rate strategy with a learning rate of 5e-5, total training steps of 30K, and a warmup period of 0.5K steps, using a batch size of 1 and increasing the gradient accumulation to 8.
Speaker-Adaptive Lip Reading Model For vision level adaptation, the training steps are set to 300 updates. We use a cosine learning rate scheduler without warmup updates. The learning rate begins at 1e-4 and gradually decreases to a minimum of 1e-5 over 5000 update periods. The LoRA configuration includes a rank of 8 and a scaling factor of 16. The target modules for LoRA adaptation are the convolutional layer, and a query, a key, and a value in the self-attention layer. For language level adaptation, the visual encoder is kept frozen to focus adaptation efforts on the language model. The LoRA configuration is applied to the attention layers of the LLM, specifically targeting the query, key, and value components. The number of training updates is set from 300 to 70.
Baselines for comparisons
Since this is the first work to explore the effectiveness of the speaker-adaptive lip reading at the wild sentence level, we set comparison methods including previous word-level speaker-adaptive methods and full fine-tuning methods, to validate the effectiveness of the proposed method.
@@
Table 3: This table compares the WER for various speaker-adaptive training strategies, detailing changes in trainable parameters for the encoder and decoder. indicates the number of total parameters.
@@
Comparsion with previous method at Vision and Language Levels Adaptation.
Table 3. provides details of experiments conducted to evaluate different training strategies in speaker adaptive lip reading. The baseline method consists of a considerable number of parameters, with 181.3 million in the encoder and 8 billion in the decoder, resulting in a Word Error Rate (WER) of 47.3%.
@@
Table 4: WER Performance ($\downarrow$) Across Different Adaptation Durations: This table shows the WER for 20 speakers, across different adaptation durations from 1 up to 45 minutes. Each duration denotes the total length of the video sample used for adaptation. This result analyzes how the model’s performance varies with each duration of adaptation data, providing crucial information for determining the minimal adaptation time.
In the Vision Level Adaptation, we compare the proposed method with V LoRA, Padding Prompt, and three types of full fine-tuning to validate the effectiveness. V LoRA and Padding Prompt added a small amount of trainable parameters (0.6M and 0.1M, respectively) but achieved improvements in WER, reducing it to 42.5% and 42.9%, respectively. The full-fine-tuning methods Finetune F, B, and F&B, obtained substantial reductions, with results of 42.8%, 40.2%, and 40.0%, respectively. While these methods adapting the back-end modules show superior performances, they require training over 170M parameters. In contrast, our proposed method achieves a WER of 41.5% by training only 0.7M parameters. It is worth noting that the proposed method outperforms the Finetune-F method, which utilizes 11.2M parameters.
In the Language Level Adaptation, we have kept the parameters of the visual encoder frozen to focus on verifying the effectiveness of adapting speaker-specific language information. We compare the proposed method, which utilizes input prompt tuning, with the LoRA-based approach. By employing the LoRA approach in the LLM for language level adaptation, we obtain a WER of 44.7%. With the proposed method, a WER of 44.1% is achieved by using only 0.04M learnable parameters. It is worth noting that despite the proposed method utilizing fewer parameters compared to the LoRA approach, it shows better performance.
In both the Vision and Language Level Adaptations, we compare the proposed method with the Finetune F&B, which shows the best performance in the Vision Level Adaptation. For a fair comparison, we also train the LLM part of the pre-trained lip-reading model adapted on Finetune F&B using our input prompt. The results are shown in the lower part of Table 3. With the Finetune F&B adaptation method, a WER of 39.6% is achieved, which is the best performance in these experiments. The proposed method achieves a marginally higher WER of 40.9% than Finetune F&B by utilizing adaptations on both vision and language levels for the target speaker. Despite showing marginally lower performance compared to Finetune F&B, the proposed method has the advantage of using only 0.74M parameters, which is less than 0.1% of the number of parameters of Finetune F&B.
@@
One of the challenges in speaker-adaptive lip reading is it is not easy to collect sufficient data for the target speaker. Therefore, we validate the effectiveness of the proposed method by using a small amount of adaptation data corresponding to 1 and 5 minutes. Moreover, in order to verify the effectiveness according to the amount of adaptation data, we additionally conduct experiments by utilizing 15, 30, and 45 minutes of adaptation data.
Table 4 illustrates the effect on target speaker lip reading performance according to variations of the amount of speaker-specific data, across 20 speakers. The baseline WER scores across 20 speakers range widely, with the lowest WER observed at 29.6 and the highest at 83.8, demonstrating significant variability in initial pre-trained model accuracy across different speakers. The overall mean WER for the baseline is calculated at 47.3.
When we apply the proposed method across 20 speakers, the average performance is increased in all settings, regardless of the amount of adaptation data. Specifically, the speaker-adaptive lip reading model trained by 1-minute speaker-specific data achieves a better average of 46.4% WER, compared to the baseline. However, the performance of some speakers is worse than the baseline. This result is different from (Kim, Kim, and Ro 2022) shows that 1 minute of adaptation data is enough to improve the performance of all speakers in word-level lip reading. In sentence-level speaker adaptive lip reading, it seems that there is a need to be over 1 minute of speaker-specific data, generally to adapt the pre-trained model to the target speaker.
By analyzing the results when using 5 minutes of speaker-specific data, the performance of the adaptive lip reading model outperforms that of the baseline for 19 out of 20 speakers. In other words, utilizing 5 minutes of speaker-specific data is enough to adapt the pre-trained model to target speakers. The adapted model achieves an average WER of 43.4%, which is a 3.9% improvement to the baseline model. Overall, employing more speaker-specific training data up to 45 minutes improves the average WER without saturation, as seen in the average WERs of 42.5%, 41.4%, and 40.9% for 15, 30, and 45 minutes, respectively.
@@ Conclusion
This study has introduced a speaker-adaptive lip reading method that enhances the adaptation of a pre-trained lip reading model to specific target speakers by integrating both visual and linguistic cues. Specifically, we adapted the visual encoder to visual information of the target speaker, such as lip appearances and speaking speed, by using a padding prompt and LoRA. Moreover, we adapted the LLM to the speaker-specific language information by utilizing the input prompt tuning. Our novel dataset, VoxLRS-SA, derived from VoxCeleb2 and LRS3, significantly has expanded the vocabulary and poses variations available for testing, addressing the limitations of previous datasets and facilitating more robust validation in realistic scenarios. The experimental results have shown the necessity of speaker-adaptive lip reading in real-world scenarios. Moreover, our method not only improves sentence-level lip reading using only a small amount of data but also outperforms existing speaker-adaptive methods. These findings suggest that incorporating both visual adaptation and language-specific characteristics of speakers can substantially benefit lip reading technologies, especially in real-world environments.

### v2

@@ Introduction
Preprocessing By following a preprocessing process (Ma et al. 2023), we crop the mouth region from talking face videos into a size of 96 pixels in both width and height. We employ RetinaFace (Deng et al. 2020) to detect facial landmarks and crop the mouth region based on these landmarks. For data augmentation (Shi et al. 2022), horizontally flipping and randomly cropping are used.
Architecture The design of our visual encoder incorporates a conformer architecture, configured with an embedding dimension of 768, 12 attention heads, a feed-forward dimension of 3072, and 12 blocks. The configuration of the transformer decoder, which is used for pre-training the visual encoder, includes an embedding dimension of 768, 8 attention heads, a feed-forward dimension of 3072, and 9 blocks. We employ the LLaMA3-8B (Touvron et al. 2023) model as the LLM for our proposed method.
Baseline Lip Reading Model We initialize the visual encoder with weights from a pre-trained lip reading model trained on the LRS3 dataset (Ma et al. 2023). The visual encoder is fine-tuned alongside a transformer decoder from scratch on the Baseline part of the VoxLRS-SA dataset. During the first stage of training, we employ a tri-stage learning rate with a peak learning rate of 1e-3, warmup steps of 10K, decay steps of 20K, and utilize 8 NVIDIA RTX 3090 GPUs with a maximum frame count of 1800 and gradient accumulation set to 4. In a further step of our experiment, we replace the transformer decoder with the frozen LLaMA3-8B model, continuing the training on the same datasets. In this second stage of training, we adopt a cosine learning rate strategy with a learning rate of 5e-5, total training steps of 30K, and a warmup period of 0.5K steps, using a batch size of 1 and increasing the gradient accumulation to 8.
Speaker-Adaptive Lip Reading Model For vision-level adaptation, a number of training steps is set to 300, and use a cosine learning rate scheduler without warmup. The learning rate begins at 1e-4 and gradually decreases to a minimum of 1e-5 over 5000 update periods. The LoRA configuration includes a rank of 8 and a scaling factor of 16. The target modules for LoRA adaptation are the convolutional layer, a query, a key, and a value in the self-attention layer. For language-level adaptation, the visual encoder is kept frozen to focus adaptation efforts on the language model. The LoRA configuration is applied to the attention layers of the LLM, specifically targeting the query, key, and value components. The number of training updates is set to 70.
@@
Table 3: This table compares the WER performance ($\downarrow$) for various speaker-adaptive training strategies, detailing changes in trainable parameters for the encoder and decoder. indicates the number of total parameters.
@@
Baselines for comparisons
Since this is the first work to explore the effectiveness of speaker-adaptive lip reading at the wild sentence level in English, we have established comparison methods including previous word-level speaker-adaptive methods and full fine-tuning methods to validate the effectiveness of the proposed method.
Comparison with the previous method at Vision and Language Levels Adaptation
Table Implementation details. provides details of experiments conducted to evaluate different training strategies in speaker adaptive lip reading. The baseline method consists of a considerable number of parameters, with 181.3 million in the encoder and 8 billion in the decoder, resulting in a Word Error Rate (WER) of 47.3%.
@@
Table 4: WER performance ($\downarrow$) across different adaptation durations: This table shows the WER for 20 speakers, across different adaptation durations from 1 up to 45 minutes. Each duration denotes the total length of the video sample used for adaptation.
In the vision-level adaptation, we compare the proposed method with V LoRA, Padding Prompt, and three types of full fine-tuning to validate the effectiveness. V LoRA and Padding Prompt added a small amount of trainable parameters (0.6M and 0.1M, respectively) but achieved improvements in WER, reducing it to 42.5% and 42.9%, respectively. The full-fine-tuning methods Finetune F, B, and F&B, obtained substantial reductions, with results of 42.8%, 40.2%, and 40.0%, respectively. While these methods adapting the back-end modules show superior performances, they require training over 170M parameters. In contrast, our proposed method achieves a WER of 41.5% by training only 0.7M parameters. It is worth noting that the proposed method outperforms the Finetune-F method, which utilizes 11.2M parameters.
In the language-level adaptation, we have kept the parameters of the visual encoder frozen to focus on verifying the effectiveness of adapting speaker-specific language information. We compare the proposed method, which utilizes input prompt tuning, with the LoRA-based approach. By employing the LoRA approach in the LLM for language level adaptation, we obtain a WER of 44.7%. With the proposed method, a WER of 44.1% is achieved by using only 0.04M learnable parameters. Despite the proposed method utilizing fewer parameters compared to the LoRA approach, it shows better performance.
In the vision-and-language level adaptation, we compare the proposed method with the Finetune F&B, which shows the best performance in the vision-level adaptation. For a fair comparison, we also apply the language-level adaptation on Finetune F&B using our input prompt. The results are shown in the lower part of Table Implementation details. With the Finetune F&B adaptation method, a WER of 39.6% is achieved, which is the best performance in these experiments. The proposed method achieves a marginally higher WER of 40.9% than Finetune F&B by utilizing adaptations on both vision and language levels for the target speaker. Despite showing marginally lower performance compared to Finetune F&B, the proposed method has the advantage of using only 0.74M parameters, which is less than 0.1% of the number of parameters of Finetune F&B.
@@
One of the challenges in speaker-adaptive lip reading is that it is not easy to collect sufficient data for the target speaker. Therefore, we validate the effectiveness of the proposed method by using a small amount of adaptation data corresponding to 1 and 5 minutes. Moreover, in order to verify the effectiveness according to the amount of adaptation data, we additionally conduct experiments by utilizing 15, 30, and 45 minutes of adaptation data.
Table 4 illustrates the effect on target speaker lip reading performance according to variations of the amount of speaker-specific data, across 20 speakers. The WER performances of a baseline across 20 speakers range widely, with the lowest WER observed at 29.6 and the highest at 83.8, demonstrating variability in pre-trained baseline lip reading model accuracy across different speakers. The overall mean WER for the baseline is calculated at 47.3.
When we apply the proposed method across 20 speakers, the average performance is increased in all settings, regardless of the amount of adaptation data. Specifically, the speaker-adaptive lip reading model trained using 1-minute speaker-specific data achieves an average of 46.2% WER, which is 1.2% better than the baseline. However, the performance of some speakers is worse than the baseline. This result is different from (Kim, Kim, and Ro 2022), which shows that 1 minute of adaptation data is enough to improve the performance of all speakers in word-level lip reading. In sentence-level speaker adaptive lip reading, it seems that there is a need to be over 1 minute of speaker-specific data, to adapt the pre-trained model to the target speaker.
By analyzing the results when using 5 minutes of speaker-specific data, the performance of the adaptive lip reading model outperforms that of the baseline for 19 out of 20 speakers. In other words, utilizing 5 minutes of speaker-specific data is enough to adapt the pre-trained model to target speakers. The adapted model achieves an average WER of 43.4%, which is a 3.9% improvement to the baseline model. Overall, employing more speaker-specific training data up to 45 minutes improves the average WER without saturation of performances, as seen in the average WERs of 42.5%, 41.4%, and 40.9% for 15, 30, and 45 minutes, respectively.
@@ Conclusion
This study has introduced a novel speaker-adaptive lip reading method that can be adapted to specific target speakers at both the vision and language levels. Specifically, we have adapted the visual encoder to the visual information of the target speaker, such as lip appearances and speaking speed, by using padding prompts and LoRA. Moreover, we have also adapted the LLM to speaker-specific language information by utilizing input prompt tuning for language-level adaptation. To validate the effectiveness of the proposed method, we have introduced a novel VoxLRS-SA dataset that expanded the vocabulary size and poses variations, compared to previous speaker-adaptive datasets in English. The experimental results have shown the necessity of speaker-adaptive lip reading in real-world scenarios. Moreover, our method not only improves sentence-level lip reading by using only a small amount of data but also outperforms existing speaker-adaptive methods. These findings suggest that incorporating both visual adaptation and language-specific characteristics of speakers can substantially benefit lip reading technologies.
Acknowledgments
This work was partly supported by two funds: the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2C2005529) and IITP grant funded by the Korea government(MSIT) (No.2020-0-00004, Development of Previsional Intelligence based on Long-Term Visual Memory Network)
Appendix A Additional Results
Selecting Optimal Weights for Lora
We have conducted eight additional experiments to assess the effectiveness of LoRA across different weight types in vision-level adaptation, as detailed in Table 5. The components $W_{c}$, $W_{q}$, $W_{k}$, and $W_{v}$ represent the LoRA weights for convolution, query, key, and value, respectively, with performance evaluated using WER. Among them, $W_{v}$ showed the most significant impact when applied LoRA to a single weight type, reducing the WER to 39.6%. Applying LoRA to combined weights, the configuration of $W_{q}$, $W_{k}$, $W_{v}$ achieved a WER of 38.5%. Extending this adaptation to include $W_{c}$, resulting in the configuration $W_{c}$, $W_{q}$, $W_{k}$, $W_{v}$, achieved a WER of 38.6%, similar to the previous configuration. Furthermore, configurations at different ranks—8, 4, and 2—have achieved WERs of 38.6%, 38.9%, and 38.4%, respectively. These results are significantly better compared to the baseline WER of 42.2%.
Failure Cases
We have evaluated failure cases, where our proposed method shows higher WER than the baseline, and noticed two general patterns: misinterpretation of context and substitution errors. Our proposed method may distort the intended context by focusing on individual speaking styles rather than the actual spoken words. For example, when the ground truth is ”but donald trump had nothing but praise for theresa may in this interview wants to see”, our proposed method predicts as ”that technology has nothing to do with the rise of making this even worse”. Also, even with the speaker-adaptive approach, there are instances where our method substitutes words with similar-sounding words. For example, the ground truth is ”and likes prime minister is people know that he is not interested in the office for what it can give him he’s interested because of public service and he thinks two terms at the end of those two terms”, and the prediction is “that line’s premise is people know that he’s not interested in the office for what he’ll give him he’s interested in the plans of public service and he says two times and the end of those two times”. The substitution of ”likes” with ”that line’s,” ”prime minister” with ”premise,” and ”he’ll” with ”he’ll give him” demonstrates that the method occasionally struggles to capture intended words.
Appendix B Limitation
While we have utilized efficient adaptation methods such as LoRA and prompt tuning, which require only minimal adjustments to the pre-trained model’s parameters, these methods still necessitate further training. Moreover, although our proposed method showed promising results when adapting the model using only 5 minutes of speaker-specific data, obtaining such data from a broad range of speakers is not easy. Therefore, as a future direction, we expect to explore methods that can effectively learn from even less data to mitigate this limitation, thus improving the adaptability and practicality of the model.
  Weight Type Rank $r$ WER(%)
Baseline - 42.2
[…]
$W_{q}$ 8 41.0
$W_{k}$ 8 40.6
$W_{v}$ 8 39.6
$W_{q}$, $W_{k}$, $W_{v}$ 8 38.5
$W_{c}$, $W_{q}$, $W_{k}$, $W_{v}$ 8 38.6
$W_{c}$, $W_{q}$, $W_{k}$, $W_{v}$ 4 38.9
$W_{c}$, $W_{q}$, $W_{k}$, $W_{v}$ 2 38.4
Table 5: Comparative analysis of WER performance ($\downarrow$) across various LoRA weight configurations and ranks in vision-level adaptation, tested on the S1 speaker from the VoxLRS-SA dataset.
Figure 2: Examples of failure cases highlighting misinterpretation of context and substitution errors in predictions